A U-shaped long-run average total cost curve can be explained by firms increasing their factory size to
(x) avoid coordination problems that occur when the factory is large.
(y) take advantage of greater specialization.
(z) avoid fixed costs.
A. (x), (y) and (z)
B. (x) and (y) only
C. (x) and (z) only
D. (y) and (z) only
E. (y) only

A business firm produces and sells specialty cakes. Last year, the firm produced 12,000 cakes and sold each cake for $20. In producing the 12,000 cakes, it incurred variable costs of $150,000 and a total cost of $183,000. Which of the following statements is (are) correct?
(x) The firm’s economic profit for the year was more than $51,250 but less than $54,375.
(y) Fixed costs amounted to $33,000 and average fixed costs are $2.75 per unit for 12,000 cakes.
(z) In producing the 12,000 specialty cakes, the firm’s average variable cost was more than $12.25 per cake and its average total cost was less than $15.75 per cake.
A. (x), (y) and (z) B. (x) and (y) only
C. (x) and (z) only D. (y) and (z) only E. (y) only

Which of the following statements is (are) correct?
(x) If marginal cost is $10 and rising, average variable cost could be either greater than or less than $10.
(y) If you know the average variable cost of 75 units then you have sufficient information to calculate the marginal cost of the 75th unit.
(z) If you know the average fixed cost of 600 units then you have sufficient information to calculate the average fixed cost of 400 units.
A. (x), (y) and (z) B. (x) and (y) only
C. (x) and (z) only D. (y) and (z) only
E. (x) only
Answer #1

Answer-1 The correct option is B.) (x) and (y) only

Long-run average total cost curves are often U-shaped because of increasing specialization of workers
at low levels of production and increasing coordination problems at high levels of production.

Answer-2 The correct option is E.) (y) only

Total Revenue= Price * Quantity

                       = 20 * 12000

                       = $2,40,000

Variable costs = $150,000

Total cost =  $183,000.

Economic Profit= Total Revenue- Total Cost

                          = 240000- 183000

                         = 57,000

Fixed cost = Total cost - Variable cost

                 = 183,000- 150,000

                 =33,000

Average fixed cost = Fixed cost/Quantity

                              = 33,000/12,000

                              = 2.75

Average variable cost = Variable cost/Quantity

                                    = 150,000/12,000

                                    = 12.5

Average total cost = Total cost/Quantity

                            = 183,000/12,000

                                    = 15.25

Answer-3 The correct option is C. )(x) and (z) only

(y) is not correct because by information of average variable cost of 75 units we cannot calculate calculate the marginal cost of the 75th unit. We need to have information of average variable cost of 74 units.